OpenStudy (clamin):

PLEASE HELP!!! MEDAL!! Simplify the expression:: 7r^2-49r/r-7

OpenStudy (danjs):

where are the parenthesis to group the numerator and denominator as typed in you have...

OpenStudy (danjs):

\[7r^2 - 49\frac{ r }{ r } -7\]

OpenStudy (danjs):

I assume you mean (7r^2 -49r) / (r-7)

OpenStudy (danjs):

\[\frac{ 7r^2-49r }{ r-7 }\]

OpenStudy (clamin):

idk..cause oon my worksheet there is no parenthesis

OpenStudy (danjs):

ok, well if it is just this 7r^2-49r/r-7 then that simplifies to 7r^2 - 49 - 7 = 7r^2 - 56

OpenStudy (clamin):

the one that you wrote was right

OpenStudy (danjs):

But if it is grouped like i did above, it is factoring the numerator, taking out a 7r, and canceling the r - 7 term \[\frac{ 7r^2-49r }{ r-7 } = \frac{ 7r(r-7) }{ r-7 } = 7r\]

OpenStudy (clamin):

@Miracrown

OpenStudy (danjs):

? you understand what happened?

OpenStudy (clamin):

no

OpenStudy (triciaal):

@clamin r^2 is r*r and 49 is 7*7 look at the numerator

OpenStudy (triciaal):